# xxx EFI-25-4

### v1

@@ 1 Introduction
In this paper, we take a fourth approach to solving the equations of motion of EFTs. This is based on the observation that in an EFT there are two types of redundancies at the level of the action. Firstly, total derivatives do not affect the equations of motion, and hence such terms can be removed from the action. Secondly, certain terms can be removed (or added) by performing perturbative field redefinitions. It is well-known that EFTs related by perturbative field redefinitions describe the same low energy physics (see e.g., [1, 34], and also [35, 36] for related statements on the invariance of the S-matrix). More recently, it was discovered by [37, 38, 39, 40] in the context of first order relativistic viscous hydrodynamics and by [41] in the context of a general class of gravitational theories that field redefinitions can be exploited to modify the character of the equations of motion and obtain a well-posed initial formulation of EFTs. In particular, field redefinitions that perturbatively shift the light fields by terms that vanish on-shell in the zeroth order theory introduce a special set of higher derivative terms into the equations. These higher derivative terms inherit the good high frequency behaviour of the zeroth order equations of motion.
@@
Finally, the regularisation scheme introduces a number of massive ghost fields. One may be concerned that despite our results on local well-posedness and a careful choice of initial data, solutions may still exhibit runaway behaviour. Indeed, due to the nonlinear interactions between the light fields and the ghost fields, the latter will inevitably get excited to some extent during the evolution. Then, energy conservation would allow for a runaway solution, since the ghost fields can carry arbitrarily negative energy while the light fields can carry arbitrarily large positive energy. However, this argument is not a proof of the inevitability of a cascading behaviour for all EFT-compatible initial data, and there may be dynamical obstructions that prevent it. In fact, recent works [42, 43, 44] and the results of this paper support this claim. We will return to this issue in the Discussion section (see the subsection ‘On the global nonlinear stability of the vacuum’).
@@ 2.1 EFT expansion
[…]
@@ 3.1 Linearisation
$||\theta_{\rm UV}-\theta_{\text{EFT}_{0}}||_{C_{t}^{0}L^{2}_{x}}\leq\frac{C\,t^{\lfloor\frac{l}{2}\rfloor+2}}{M^{l+1}}\,,$ (3.7)
where $C$ is a constant, $t\leq\tilde{C}\,M^{\lambda}$ for some other constant $\tilde{C}$, and $20\leq\lambda<$ 2. The precise definition of the $C_{t}^{0}L^{2}_{x}$ norm can found in Section 4. In Section 5 we perform some numerical experiments to verify the expected accuracy (3.7) of the EFT0 solutions.
@@ 3.2 Second order EFTs
Considering plane wave solutions $\delta\theta=\Theta\,e^{\text{i}k\cdot x}$ with $\Theta=\text{const.}$., we obtain the dispersion relation
@@ 3.4 Initial data
$(\partial_{t}^{3}\theta)\Big{|}_{t=0}=$ $\,\frac{1}{[M^{2}-2(\partial_{x}\theta_{0})^{2}+6\,\theta_{1}^{2}]^{3}}\Big{[}16\,\partial_{x}\theta_{1}\,\partial_{x}\theta_{0}\,\partial_{x}^{2}\theta_{0}\Big{(}-6\,\theta_{1}^{4}+6\,\big{(}\partial_{x}\theta_{0})^{4}-3\,M^{2}\,(\partial_{x}\theta_{0})^{2}$ (3.42)
$\hskip 142.26378pt+\theta_{1}^{2}(96\,(\partial_{x}\theta_{0})^{2}-13\,M^{2})\big{)}\Big{)}$
$+\partial_{x}^{2}\theta_{1}\left(6\,\theta_{1}^{2}-2\,(\partial_{x}\theta_{0})^{2}+M^{2}\right)\big{(}12\,\theta_{1}^{4}+12\,(\partial_{x}\theta_{0})^{4}-8\,M^{2}\,(\partial_{x}\theta_{0})^{2}+M^{4}$
@@
$-8\,\theta_{1}\,(\partial_{x}^{2}\theta_{0})^{2}\big{(}48(\partial_{x}\theta_{0})^{2}-6\,M^{2}\,(\partial_{x}\theta_{0})^{2}+M^{4}+2\,\theta_{1}^{2}(24(\partial_{x}\theta_{0})^{2}+M^{2})\big{)}$
@@ 4.1 Notations and definitions of norms
$||u||_{H^{s}}=\left(\sum\limits_{k\leq s}\int{\rm d}x\,|\partial_{x}^{k}u|^{2}\right)^{1/2}.$ (4.1)
@@
$||u||_{C^{0}_{t}H^{s}_{x}}=\sup_{t\in[0,T]}||u(t,\cdot)||_{H^{s}}\,,$ (4.2)
which can be thought of as the upper envelope of the function $||u||_{H^{s}_{x}}(t)$.
@@ 5 Numerical experiments
[…]
@@ 5.2 Long time behaviour of the errors
On the other hand, the regularisation scheme provides a single non-linear equation corresponding to any EFT truncated at some given order in the derivative expansion. [^7] In our numerical scheme we solve a system of second order wave equations rather than a single higher order nonlinear equation. However, in this formulation there is only one equation that contains a truncation error, the rest of the equations are simply the defining equations of the auxiliary fields and hence they have vanishing error. If the parameters in the field redefinition are chosen such that the masses of the auxiliary fields are positive, we expect that (in the regime of validity of EFT) both the solutions of the regularised EFT and the truncation errors in the equations of motion are bounded (for sufficiently long times). Therefore, the argument in the previous paragraph implies that the error with respect to the full solution is expected to grow linearly in time.
@@ 5.4 Breakdown of the EFT
$-\frac{1}{\mu^{2}}\Box(\Box\theta)+A^{ab}\partial_{a}\partial_{b}\theta+B^{a}\partial_{a}\theta+C\,\theta=0\,,$ (5.5)
@@ 5.5 Gaussian initial data
[…]
@@ 6 Discussion
We carried out simple numerical experiments using the initial data (5.1) for the massless scalar field $\theta$. Consistent initial data for the massive UV field $\rho$ can obtained from the EFT expansion and using the reduction of order procedure. Similarly, initial data for the auxiliary fields (higher derivatives of $\theta$) in the EFT2 and EFT4 is obtained by order reduction. Notably, the initial data (5.1) has an amplitude of $O(1)$ hence the non-linearities in the equations play an important role on the initial slice already. Varying the mass scale $M$ while keeping the initial data fixed allows us to interpolate between the regime of validity of EFT and a regime where the EFT expansion breaks down. In the regime of validity of EFT, we see a clear imporvement in the accuracy of the solutions provided by EFT4 compared to the solutions provided by EFT1 and EFT2. The latter two EFTs provide equivalent descriptions in agreement with the expectations. We have also checked that the errors in the solutions constructed with the EFTs scale with the expected powers of the UV mass scale $M$. Furthermore, since we solve the evolution equations of these EFTs fully non-linearly, the late time error of these EFT solutions grows linearly in time, which is optimal. This should be contrasted with the errors of the solutions constructed with the linearisation scheme, which grow polynomially in time, with a power that depends on the truncation order of the EFT (3.7).
[…]
@@
It is interesting to note that the global behaviour of solutions is expected to be better when the ghost fields are massive (which is the case in the regularised EFTs discussed in this paper), as also recently observed by [44]. This is because massive fields generally exhibit faster decay than massless fields. The improved stability can also be understood heuristically by noticing that the mass term creates a well in the potential, see Fig. 11. The field $\Phi$ (which here is a representative of a ghost field, i.e., the analogue of a higher derivative of the light field $\theta$ in the EFTs discussed in this paper) has to travel a distance of $O(M)$ in field space to overcome the potential barrier and exhibit runaway behaviour.
@@ Acknowledgements
We would like to thank Ramiro Cayuso, Katy Clough, Aaron Held, Luis Lehner, Eugene Lim and Bob Wald for discussions. PF and ADK are supported by the STFC Consolidated Grant ST/X000931/1. SY acknowledges the support from the China Scholarship Council.
@@ Appendix A Consistency checks
In this section we carry out some consistency checks of our numerical simulations. As discussed in Section 4, the UV theory and the various EFTs admit conserved charges. To check the accuracy of our numerical simulations, we monitored the conservation of the energy and the charge associated with the shift symmetry throughout the evolution. Here we present results for the initial data (5.1) and $M=100$. In this setup, both the energy ${\cal E}$ and the charge ${\cal Q}$ are positive in each of the theories considered. In Fig. 13 we demonstrate that ${\cal E}$ and ${\cal Q}$ remain constant (within numerical accuracy) in time throughout our simulations. Note that for each theory, the expressions for ${\cal E}$ and ${\cal Q}$ are different, and hence these plots merely serve the purpose of checking the consistency of our simulations for each theory.
@@
Therefore, if the constraints are initially satisfied then they will continue to hold at all times. Of course, in numerical simulations the finite resolution will inevitably introduce some constraint violations. To quantify the accuracy to which we solve the original higher derivative equations, we compute the relative constraint violation $||{\cal C}_{A}||_{L^{2}}/||\theta_{A}||_{L^{2}}$ for our auxiliary variables. In Fig. 14 we show the relative errors for these variables in $\text{EFT}_{4}$ with $M=100$ and initial data (5.1), confirming that the constraint violations are negligibly small.

### v2

@@ 1 Introduction
In this paper, we take a fourth approach to solving the equations of motion of EFTs. This is based on the observation that in an EFT there are two types of redundancies at the level of the action. Firstly, total derivatives do not affect the equations of motion, and hence such terms can be removed from the action. Secondly, certain terms can be removed (or added) by performing perturbative field redefinitions. It is well-known that EFTs related by perturbative field redefinitions describe the same low energy physics (see e.g., [1, 34], and also [35, 36, 37, 38] for related statements on the invariance of the S-matrix). More recently, it was discovered by [39, 40, 41, 42] in the context of first order relativistic viscous hydrodynamics and by [43] in the context of a general class of gravitational theories that field redefinitions can be exploited to modify the character of the equations of motion and obtain a well-posed initial formulation of EFTs. In particular, field redefinitions that perturbatively shift the light fields by terms that vanish on-shell in the zeroth order theory introduce a special set of higher derivative terms into the equations. These higher derivative terms inherit the good high frequency behaviour of the zeroth order equations of motion.
@@
Finally, the regularisation scheme introduces a number of massive ghost fields. One may be concerned that despite our results on local well-posedness and a careful choice of initial data, solutions may still exhibit runaway behaviour. Indeed, due to the nonlinear interactions between the light fields and the ghost fields, the latter will inevitably get excited to some extent during the evolution. Then, energy conservation would allow for a runaway solution, since the ghost fields can carry arbitrarily negative energy while the light fields can carry arbitrarily large positive energy. However, this argument is not a proof of the inevitability of a cascading behaviour for all EFT-compatible initial data, and there may be dynamical obstructions that prevent it. In fact, recent works [44, 45, 46, 47] and the results of this paper support this claim. We will return to this issue in the Discussion section (see the subsection ‘On the global nonlinear stability of the vacuum’).
@@ 2.1 EFT expansion
$\frac{S}{v^{2}}\approx-\int{\rm d}^{d}x\bigg\{$ $-\frac{1}{2}\,\mathcal{F}_{2}-\frac{1}{2M^{2}}\,\mathcal{F}_{2}^{2}+\frac{1}{2M^{4}}(\partial_{a}\mathcal{F}_{2})(\partial^{a}\mathcal{F}_{2})$ (2.22)
@@ 3.1 Linearisation
$\theta_{\rm UV}-\theta_{\text{EFT}_{0}}||_{C_{t}^{0}L^{2}_{x}}\leq\frac{C\,t^{\lfloor\frac{l}{2}\rfloor+2}}{M^{l+1}}\,$, (3.7)
where $C$ is a constant, $t\leq\tilde{C}\,M^{\lambda}$ for some other constant $\tilde{C}$, and $0\leq\lambda<2$. The precise definition of the $C_{t}^{0}L^{2}_{x}$ norm can found in Section 4. In Section 5 we perform some numerical experiments to verify the expected accuracy (3.7) of the EFT0 solutions.
@@ 3.2 Second order EFTs
Considering plane wave solutions $\delta\theta=\Theta\,e^{\text{i}k\cdot x}$ with $\Theta=\text{const.}$, we obtain the dispersion relation
@@ 3.4 Initial data
[…]
@@
[…]
@@ 4.1 Notations and definitions of norms
$u||_{H^{s}}=\left(\sum\limits_{k\leq s}\int{\rm d}x\,|\partial_{x}^{k}u|^{2}\right)^{1/2}.$ (4.1)
@@
[…]
which can be thought of as the upper envelope of the function $u||_{H^{s}_{x}}(t)$.
@@ 5 Numerical experiments
In this Section we carry out numerical experiments to explore the dynamics of the UV theory and the various EFTs for different classes of initial conditions and values of the mass scale $M$. For simplicity, we carry out the simulations in (1+1)-dimensional Minkowski space with a compact spatial domain $x\in[-1,\,1]$ and periodic boundary conditions. Therefore, the length $L$ of the spatial domain is $L=2$. These simulations are simple enough so that Mathematica’s NDSolve function suffices; we use fourth order finite differences and a fourth order Runge-Kutta time integrator unless otherwise stated.
@@
Figure 2: The spectral density of the fields plotted in Fig. 1: $\rho$ (top left), $\theta$ (top right), $\theta^{(0,1)}$ (bottom left) and $\theta^{(0,2)}$ (bottom right).
To obtain further information about the Fourier spectrum of the solution, we numerically computed the Fourier transform of the solutions for $\theta_{\rm UV}$, $\rho_{UV}$ in the UV theory and $\theta^{(0,1)}$, $\theta^{(0,2)}$ in EFT4. We show the spectral density of these fields in Fig. 2. It is clear that $\theta_{\rm UV}$ and $\rho_{\rm UV}$ are dominated by low frequency modes: the amplitudes of modes with frequencies $\gtrsim M$ are highly suppressed (although there is a small peak in $|\hat{\rho}_{\rm UV}|^{2}$ at around $\omega\sim M$ as shown on the inset of the top left panel). This behaviour of the UV solution is also captured well by the solution to EFT4. The spectral density for $\theta_{\text{EFT}_{4}}$ is indistinguishable from that of $\theta_{\rm UV}$ on the scale of the plots, hence we only include one of the in Fig. 2 in order to avoid repetition. On the other hand, the regularising terms have a small but noticeable effect on derivatives of $\theta$: in the spectral densities of $\theta^{(0,1)}$, $\theta^{(0,2)}$ in EFT4 (displayed in the bottom left and right panels of Fig. 2) have small peaks around frequencies equal to the masses of the auxiliary massive fields. Nevertheless, the amplitudes of these high frequency modes are several orders of magnitude smaller than the amplitudes of the relevant low frequency modes and these deviations lie beyond the expected accuracy of EFT4. These results suggest that the regularising terms do not affect the solution below the cutoff frequency.
@@ 5.2 Long time behaviour of the errors
On the other hand, the regularisation scheme provides a single non-linear equation corresponding to any EFT truncated at some given order in the derivative expansion. [^7] In our numerical scheme we solve a system of second order wave equations rather than a single higher order nonlinear equation. However, in this formulation there is only one equation that contains a truncation error, the rest of the equations are simply the defining equations of the auxiliary fields and hence they have vanishing error. Equation (3.14) of EFT2 contains a truncation error of $O(M^{-4})$ compared to the UV equations of motion and, similarly, equation (3.27) of EFT4 has an error $O(M^{-6})$. Assuming smooth initial data and that the parameters in the field redefinition are chosen such that the masses of the auxiliary fields are positive, we expect that (in the regime of validity of EFT) both the EFT and UV solutions and their derivatives will be uniformly bounded for sufficiently long times. Then the truncation errors in the regularised EFT equations of motion will be bounded. Based on this (and on the argument of the previous paragraph), it follows that at time $T$ the solutions of (3.14) and (3.27) will have errors $O(T/M^{4})$ and $O(T/M^{6})$, respectively. Therefore, the errors of the solutions to the regularised EFTs (measured from the corresponding UV solutions) are expected to grow linearly in time.
@@ 5.4 Breakdown of the EFT
$\frac{1}{\mu^{2}}\Box(\Box\theta)+A^{ab}\partial_{a}\partial_{b}\theta+B^{a}\partial_{a}\theta+C\,\theta=0\,$, (5.5)
@@
Figure 11: Left: The spatial Fourier transform of $\theta^{(0,1)}_{{\rm EFT}_{2}}$ computed before (at $t/L=0$) and after (at $t/L=9$) the breakdown of EFT2, showing a growth of the amplitudes of high wavenumber modes. Right: The evolution of the amplitude of the $k=0.94M$ spatial Fourier mode of $\theta^{(0,1)}_{{\rm EFT}_{2}}$. This Fourier mode has fairly small amplitude at the early stages of the evolution but exhibits rapid growth at late times.
It is also interesting to analyse the breakdown of EFT2 in Fourier space. We computed the spatial Fourier transform of $\theta^{(0,1)}_{{\rm EFT}_{2}}$ and followed the time evolution of the various Fourier modes. The results are shown on Fig. 11. On the left panel of Fig. 11 we show the amplitudes of the Fourier modes on the initial data slice (blue dots) and after the breakdown at $t/L=9$ (red dots). The initial data already has a mode $k\sim M$ with amplitude $O(1)$ (i.e. the same amplitude as other low wavenumber modes) which already indicates that EFT may not be applicable. At later times we can see significant growth of the amplitudes of modes with $k\gtrsim M$, indicating a cascade towards wavenumbers beyond the EFT cutoff. On the right panel of Fig. 11 we track the time evolution of the mode with wavenumber $k=0.94M$. This mode is already present in the initial data and exhibits small amplitude oscillations until $t\sim 4L$. Afterwards, however, the amplitude of this mode grows rapidly and becomes much larger than the amplitudes of low wavenumber modes, which is a clear indication of the breakdown of the effective field theory description.
@@ 5.5 Gaussian initial data
We fix $M=200$ and choose $a=b=40$, $c\in\{0,\,1\}$ such that the initial data has $O(1)$ amplitude and $|\partial\theta|\sim M/5$ initially. This means that initially the EFTs are quite strongly coupled and the nonlinearites are significant. For this reason, we use an implicit time integrator to evolve this initial data. Choosing larger amplitudes or steeper initial gradients while keeping $M$ fixed would lead to the blow-up of EFT solutions after a fairly short amount of time, similarly to what we have seen in Section 5.4.
@@ 6 Discussion
We carried out simple numerical experiments using the initial data (5.1) for the massless scalar field $\theta$. Consistent initial data for the massive UV field $\rho$ can obtained from the EFT expansion and using the reduction of order procedure. Similarly, initial data for the auxiliary fields (higher derivatives of $\theta$) in the EFT2 and EFT4 is obtained by order reduction. Notably, the initial data (5.1) has an amplitude of $O(1)$ hence the non-linearities in the equations play an important role on the initial slice already. Varying the mass scale $M$ while keeping the initial data fixed allows us to interpolate between the regime of validity of EFT and a regime where the EFT expansion breaks down. In the regime of validity of EFT, we see a clear improvement in the accuracy of the solutions provided by EFT4 compared to the solutions provided by EFT1 and EFT2. The latter two EFTs provide equivalent descriptions in agreement with the expectations. We have also checked that the errors in the solutions constructed with the EFTs scale with the expected powers of the UV mass scale $M$. Furthermore, since we solve the evolution equations of these EFTs fully non-linearly, the late time error of these EFT solutions grows linearly in time, which is optimal. This should be contrasted with the errors of the solutions constructed with the linearisation scheme, which grow polynomially in time, with a power that depends on the truncation order of the EFT (3.7).
In the regularised EFTs (i.e., in EFT2 and EFT4), one has the freedom to choose the parameters appearing in the field redefinitions, (3.12) and (3.25) respectively. However, there is a restriction on these parameters imposed by the requirement that the theories linearised around the trivial solution $\theta=0$ [^9] Using the shift symmetry, there is no loss of generality in choosing this particular constant solution. have no tachyonic modes. Such a choice is always possible and does not require fine-tuning. Once a choice of these parameters is fixed (together with initial data compatible with the EFT regime), the solutions obtained with the regularisation scheme remain bounded within the computational domain. Even though the regularisation scheme introduces high frequency oscillations $\omega\gtrsim M$ into the solution, we find that the amplitudes of the high frequency modes are highly suppressed. This suggests that the regularised EFT solutions satisfy the uniform boundedness requirements of [15] and the $M\to\infty$ limit may be well-defined. Furthermore, we have confirmed that the solution for the light field $\theta$ provided by the regularised EFTs is not sensitive to the specific choices of the free parameters (and hence masses) in the field redefinitions.
@@
It is interesting to note that the global behaviour of solutions is expected to be better when the ghost fields are massive (which is the case in the regularised EFTs discussed in this paper), this was shown in rigorous mathematical works [58, 59] and also recently observed by [44, 47]. This is because massive fields generally exhibit faster decay than massless fields. The improved stability can also be understood heuristically by noticing that the mass term creates a well in the potential, see Fig. 13. The field $\Phi$ (which here is a representative of a ghost field, i.e., the analogue of a higher derivative of the light field $\theta$ in the EFTs discussed in this paper) has to travel a distance of $O(M)$ in field space to overcome the potential barrier and exhibit runaway behaviour.
@@ Acknowledgements
We would like to thank Ramiro Cayuso, Katy Clough, Aaron Held, Luis Lehner, Eugene Lim, Harvey Reall and Bob Wald for discussions. PF and ADK are supported by the STFC Consolidated Grant ST/X000931/1. SY acknowledges the support from the China Scholarship Council.
@@ Appendix A Consistency checks
[…]
@@
[…]